OpenStudy (anonymous):

Factor: 2xy - xz + 2wy - wz Which of the following is one of the factors? (x + w) (x - y) (2x + y)

OpenStudy (anonymous):

x + w rearrange 2xy + 2wy -xz -wz = 2y(x+w) -z(x+w) = (x+w)(2y-z)

OpenStudy (anonymous):

thanks do you think you can help me with sqrt 32x^3y^8 ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

32^(1/2) x^(3/2) y^4 - take each part to the power 1/2

OpenStudy (anonymous):

soo the answer is??? sorry but i have noo idea how to do that!!

OpenStudy (anonymous):

that is the answer it cant be simplified any further the square root is the same as the number or letter to the power 1/2 sqrt4 = 4 ^ 1/2 = 2 sqrt x^3 = x^ (3 *1/2) = x^(3/2) sqrt x^8 = x^(8 * 1/2) = x^4

OpenStudy (anonymous):

you could use calculator to find sqrt 32 and write that if you like so answer could be written as 5.6569 x^(3/2) y^4